# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ VII Time supersolids: CES state
Figure 4: (a) Normalized density-density correlation function $G(x,-x,t)$ for $x=10$ (blue), $x=20$ (red), and $x=30$ (green) for the simulation in Fig. 3. Black line is a quadratic polynomial fit to the amplitude of the oscillations, $A(t)=at^{2}+bt+c$. (b) Zoom of (a). (c)-(d) Same as upper row but now replacing quantum fluctuations by particle-number fluctuations. Specifically, we take as initial condition in the Truncated Wigner method $\Psi(x,0)=\sqrt{1+\delta n}\,e^{ivx}$, with $\delta n$ a Gaussian random variable with zero mean and $\sqrt{\braket{\delta n^{2}}}=0.001$.
We now go beyond mean-field and explicitly compute the dynamics of the quantum fluctuations by using the Truncated Wigner method Sinatra et al. (2002); Carusotto et al. (2008); all technical details behind the numerical simulations and the specific implementation of the Truncated Wigner method can be found in Refs. de Nova et al. (2016); de Nova and Sols (2022, 2023). As initial quantum state, we take the $T=0$ ground state in the comoving frame of the condensate. Regarding the observables of interest, we focus on computing the normalized density-density correlation function
@@
with $\delta\hat{n}(x,t)\equiv\hat{n}(x,t)-\braket{\hat{n}(x,t)}$ the density fluctuations and $\xi_{0}$ the healing length associated to the initial density $n_{0}$; we have momentarily restored dimensions in this expression for the sake of clarity. In the laboratory, density-density correlations can be measured through in situ imaging after averaging over ensembles of repetitions of the experiment Steinhauer (2016); de Nova et al. (2019); Kolobov et al. (2021).
After some transient, and once in the CES state, the correlations exhibit a seemingly periodic behavior, as shown in Fig. 3, where sharp features arise due to the synchronized emission of shock waves/solitons into the upstream/downstream region. Further insight is obtained in upper row of Fig. 4, where we plot the time dependence of the correlation function $G(x,-x,t)$ between symmetric upstream and downstream points for different values of $x$. We observe that the correlations are indeed quasi-periodic, displaying periodic oscillations whose amplitude grows in time.
@@
Figure 5: Normalized density-density correlation function $G(x,-x,t)$ for $x=10$ (left column), $x=20$ (center column), and $x=30$ (right column). Solid blue is the numerical result from a Truncated Wigner simulation and dashed red is the contribution from the temporal FNG mode, $G(x,x^{\prime},t)=r_{t}(x,t)r_{t}(x^{\prime},t)A(t)$, where $A(t)$ is numerically fitted as in Fig. 4 and $r_{t}(x,t)=\partial_{t}n(x,t)$ is computed from the time derivative of the mean-field density. Upper row: $Z=1$ and $v=0.8$. Lower row: $Z=1$ and $v=0.7$.
This behavior can be understood using the spontaneous Floquet theory developed in Sec. VI. After expanding the field operator around the CES wavefunction, we find that the density fluctuations read at linear order in the field fluctuations $\hat{\varphi}(x,t)$ as
@@
$\delta\hat{n}(x,t)$ $=$ $\hat{X}^{t}(t)r_{t}(x,t)+\hat{P}^{N}r_{N}(x,t)+\hat{P}^{F}r_{F}(x,t)$
$+$ $\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}r_{\varepsilon,\nu}(x,t)e^{-i\varepsilon t}+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}r^{*}_{\varepsilon,\nu}(x,t)e^{i\varepsilon t},$
$r_{n}(x,t)=-iz^{\dagger}_{\theta}(x,t)\sigma_{z}z_{n}(x,t)$ being the density amplitude associated to the spinor $z_{n}$. In the case of the Goldstone-Gibbs modes, the density amplitude is just
$r_{a}(x,t)=\partial_{a}|\Psi_{0}(x,t)|^{2}=\partial_{a}n(x,t),$ (172)
where $n(x,t)$ is the mean-field density, which is approximately equal to the ensemble-averaged density after neglecting higher-order corrections in the field fluctuations, $\braket{\hat{n}(x,t)}\simeq n(x,t)$.
@@
$G(x,x^{\prime},t)\simeq r_{t}(x,t)r_{t}(x^{\prime},t)A(t),~{}A(t)\equiv\frac{\braket{(\hat{X}^{t}(t))^{2}}}{n_{0}\xi^{-1}_{0}}.$ (173)
@@
Due to the ballistic evolution of $\hat{X}^{t}(t)$, we expect a quadratic polynomial dependence for $A(t)$, $A(t)=at^{2}+bt+c$. This fits well the observed growth, black lines in upper row of Fig. 4. A more detailed comparison is presented in Fig. 5, where we depict together the numerical results for $G(x,x^{\prime},t)$ from the Truncated Wigner method (solid blue), and the theoretical prediction of Eq. (173), combining the previous quadratic fit of $A(t)$ with a computation of $r_{t}(x,t)=\partial_{t}n(x,t)$ from the time derivative of the mean-field density (dashed red). A remarkable agreement is observed, confirming the ballistic motion of the quantum amplitude of the temporal FNG mode as the mechanism responsible for the quasi-periodic correlations.
@@
We note that the coefficients $a,b,c$ determining the amplitude $A(t)$ are given in terms of expectation values which are quadratic in the amplitudes $\hat{X}^{t},\hat{P}^{N},\hat{P}^{F}$, evaluated in the initial quantum state. Thus, our specific time-dependent scheme eliminates any possible ambiguity in the definition of the quantum state of the system Ribeiro et al. (2022). Nevertheless, fluctuations of the temporal FNG can be also induced by run-to-run experimental variations such as shot noise in the initial particle number. This is shown in lower row of Fig. 4, where we display the results of solely including fluctuations in the particle number of the initial condition. The observed correlation patterns are the same as in the purely quantum case (upper row), further revealing their monomode origin since particle-number fluctuations are structureless de Nova and Sols (2023).
@@
Based on these results, we propose an experimental protocol to observe the temporal FNG mode using standard techniques in analogue gravity setups Steinhauer (2016); de Nova et al. (2019); Kolobov et al. (2021). We consider an elongated quasi-1D condensate that is essentially homogeneous far from the edges, along which a localized obstacle is swept with velocity $v$. By Galilean invariance, this is equivalent to launching the condensate against the obstacle with the same velocity. For values of $v$ above the critical velocity, a CES state will be reached. Another possibility is to confine the condensate in a long ring Eckel et al. (2018) and rotate a localized potential. In both cases, at long times once in the CES regime, when the temporal FNG mode dominates the Bogoliubov dynamics, high-resolution imaging can be then used to measure the time evolution of the density profile and the density-density correlations, from where the FNG mode can be extracted as in Fig. 5.
@@
We note that the protocol used here to reach the CES state is fully deterministic at the mean-field level, which imposes the value of the global time origin $t_{0}$ of the spontaneous Floquet state. Thus, even though there is a temporal FNG mode, technically speaking the symmetry breaking of time translation invariance is not spontaneous. The reason is that our scheme fixes a $t=0$ origin in the GP equation. Regardless, at late times, due to interactions, the system enters the CES self-oscillating regime where it forgets about the initial condition and exhibits all the features of a spontaneous Floquet state. We can qualitatively understand this behavior by drawing an analogy with a conventional spatial crystal growing in the $z$-direction on top of a rigid substrate, which fixes a $z=0$ plane that breaks continuous spatial translation symmetry. In the bulk of the grown crystal, however, the system becomes insensitive to the boundary and recovers the usual properties of a crystalline structure.
@@
Nevertheless, the phase locking of $\phi_{0}$ actually helps to measure the FNG mode since, if $\phi_{0}$ were spontaneously chosen in each realization of the experiment, then $G(x,x^{\prime},t)$ would represent an angular average over $\phi_{0}\in[0,2\pi)$, therefore suppressing its time dependence. In that case, one would need to resort to the complete two-time correlation function
@@
which requires the use of non-destructive imaging in order to measure the density at two consecutive times $t>t^{\prime}$, as in Ref. Smits et al. (2018). This caveat is not needed for the measurement of correlations under the usual $U(1)$ symmetry breaking, where each realization of the experiment is insensitive to the global condensate phase because observables always commute with the total particle number, the generator of global phase transformations.
VIII Discussion and perspectives
@@
We proceed to critically discuss the results of this work and how they inscribe within the literature, outlining in the process future avenues. We have examined in Sec. III how the general results from Goldstone theorem applied to variational descriptions, Sec. II, can be translated to many-body systems when several symmetries are simultaneously broken by the quantum state, focusing on the specific case of a BEC close to zero temperature for illustrative purposes. We have found that a rather complete description of the problem is provided in terms of the generalized Gibbs ensemble, which treats all broken symmetries on equal grounds. The Lagrange multipliers associated to each conserved charge behave as generalized velocities along the orbits of the symmetry transformations, extending the well-known result that the chemical potential is the velocity of the phase. We have also proven fundamental results in Thermodynamics for the symmetry-broken state, such as the first principle or the Gibbs-Duhem relation.
@@
The Nambu-Goldstone modes associated to each spontaneously broken symmetry emerge as zero-energy modes of the excitation spectrum. Each NG mode is paired with a Gibbs mode, describing the fluctuations of the quantum state with respect to the associated conserved charge. These fluctuations arise because a symmetry-broken state cannot be an eigenstate of the charge operator that generates the corresponding broken-symmetry transformation. The orthogonalization of the Goldstone-Gibbs sector involves the Berry-Gibbs connection, a Berry connection whose variables are the conserved charges, which are not intrinsic parameters of the Hamiltonian. In contrast to the standard case, the resulting Berry-Gibbs curvature is not invariant under generalized gauge transformations, which expands the usual notion of phase transformations to include more general symmetry transformations. When extended to the complete manifold that includes the continuous parameters of the spontaneously broken symmetries, the Berry-Gibbs curvature gives rise to a symplectic structure that allows to quantize the amplitudes of the Nambu-Goldstone/Gibbs modes as conjugate coordinate/momentum operators. This analogy is translated to the dynamics, where the momenta are time-independent as they represent the quantum fluctuations of the corresponding conserved charges, while the coordinates display a ballistic motion with velocities proportional to the momenta. The physical intuition behind this dynamics is very simple: since there is no restoring force along the orbits generated by the broken-symmetry transformations, fluctuations in those directions, spanned by the NG modes, lead to unbounded motion with constant velocity. Finally, we show that the Goldstone-Gibbs sector can be diagonalized, leading to a new set of independent NG modes. Remarkably, this diagonalization only involves thermodynamic considerations, and describes how the original NG modes hybridize within the different branches of the spectrum in the low-frequency limit.
@@
The coordinate-momentum analogy between the Nambu-Goldstone and Gibbs amplitudes was originally established in Ref. Lewenstein and You (1996) in the context of the $U(1)$-symmetry breaking and the phase diffusion of a condensate. This approach was followed in Ref. Dziarmaga (2004) to also quantize the position of a soliton and study its quantum diffusion, inspiring a whole line of research in quantum solitons Sacha et al. (2009); Gangardt and Kamenev (2010); Wadkin-Snaith and Gangardt (2012); Walczak and Anglin (2011); Marchukov et al. (2020). A study of the quantization of arbitrary NG modes in condensates, including internal symmetries such as spin, was provided in Ref. Takahashi and Nitta (2015). The present work further extends those results by providing a general and systematic framework with the help of the generalized Gibbs ensemble and the Berry-Gibbs connection, which in addition allows to understand the physics of the Nambu-Goldstone and Gibbs modes at a more fundamental level. On the other hand, the emergence of a geometrical description and a symplectic structure in variational formulations is well-known in the literature and can be traced back to Ref. Kramer and Saraceno (2005) (see Ref. Kramer (2008) for a more recent review), where the manifold is generated there by the variational parameters of the ansatz and their velocities, leading to a dequantization of the problem. In contrast, the manifold variables in our description are the continuous parameters of the spontaneously broken symmetries and their associated charges, instead of variational parameters, and the symplectic structure conversely leads to a quantization of the Goldstone-Gibbs amplitudes. To the best of our knowledge, the thermodynamical and geometrical framework developed in Sec. III for the description of spontaneous symmetry breaking and the quantization of the Goldstone-Gibbs modes is original from this work and is not present in the literature.
@@
As a particular application of this formalism, we have studied in Sec. IV the case of a cnoidal wave in a superfluid, which spontaneously breaks $U(1)$ and spatial translation symmetry, extending the work of Refs. Dziarmaga (2004); Martone et al. (2021) into the generalized Gibbs ensemble. We have explicitly checked the consistency and validity of the theoretical predictions of Sec. III.
@@
The results of Secs. III, IV are thus of interest for the study of states which simultaneously break several symmetries, such as supersolids, quantum droplets, spinor and rotating condensates, or quantum solitons. They can also be easily translated to magnonic condensates Mäkinen et al. (2023), nonlinear optical fibers Drummond and Hillery (2014), and quantum fluids of light Carusotto and Ciuti (2013), as they obey nonlinear equations of motion similar to the Gross-Pitaevskii equation. Indeed, the generality of the formalism, based on a Lagrangian description, allows for a straightforward adaptation to arbitrary many-body systems, including fermions. For instance, within the $\nu=0$ quantum Hall phase diagram in graphene, states with two spontaneously broken symmetries were predicted within a Hartree-Fock description de Nova and Zapata (2017), of which some experimental evidence was recently reported Coissard et al. (2022). Indeed, Ref. de Nova and Zapata (2017) established the correspondence between the BdG and TDHFA equations, and translated Goldstone theorem into the specific context of the Hartree-Fock equations. Therefore, the BdG/TDHFA correspondence provides a direct link through which our results can be exported.
@@
Another interesting perspective is provided by the geometry of the extended Berry-Gibbs connection. In this work, we have assumed that all the generators of the spontaneously broken symmetries commute between themselves. This leads, after a generalized gauge transformation, to a “flat” symplectic curvature $F_{ab}=-\Omega_{ab}$. Actually, a generalized gauge transformation is nothing else than a change of coordinates in the extended Berry-Gibbs manifold. However, if the generators of the spontaneously broken symmetries form a non-abelian Lie algebra, this no longer holds, opening the door for deeper geometrical implications, including nontrivial topological aspects. The exploration of this promising avenue is left for future work.
@@
The formalism developed for stationary states in Secs. II-III is adapted to Floquet states in Secs. V, VI. Specifically, from the general results of Sec. II, we prove that the usual Nambu-Goldstone modes with zero energy are now translated into Floquet-Nambu-Goldstone modes with zero quasi-energy. Moreover, since spontaneous Floquet states break continuous time-translational symmetry, a genuine temporal FNG mode emerges. This is a distinctive signature of a spontaneous Floquet state, completely absent in conventional Floquet systems.
@@
In order to correctly describe the dynamics of the FNG modes, we develop the $(t,\phi)$ formalism, which goes beyond the usual $(t,t^{\prime})$ formalism. In combination with the generalized Gibbs ensemble, we provide a thermodynamical description of Floquet states analogous to that of stationary states, which we refer to as Floquet thermodynamics. This is possible because spontaneous Floquet states also have a well-defined energy, which enables the existence of a conserved Floquet charge. The Floquet charge is another hallmark of a spontaneous Floquet state, which adds to the presence of a temporal Floquet-Nambu-Goldstone mode. In fact, both features are the sides of the same coin, since they arise, via Goldstone and Noether theorems, from the continuous time-translation symmetry of the underlying time-independent Hamiltonian.
@@
We stress that the use of the generalized Gibbs ensemble in combination with the $(t,\phi)$ formalism is essential for a correct and complete description of the problem. Interestingly, within this framework, a remarkable correspondence between stationary states and spontaneous Floquet states can be established, consisting of replacing stationary states by periodic states, frequencies by quasi-frequencies, and conserved charges by their angular average.
@@
Finally, we apply our formalism to a specific realization of spontaneous Floquet state, the CES state, which spontaneously breaks $U(1)$ and continuous time-translation symmetry, representing a time supersolid. We propose an experimental scheme, based on standard analogue gravity techniques Steinhauer (2016); de Nova et al. (2019); Kolobov et al. (2021), to observe the temporal FNG mode, which does couple to the density in contrast to the FNG mode of the phase. In particular, we prove that the density-density correlations are dominated at long times by the ballistic growth of the temporal FNG mode, from where it can be extracted. Numerical results from a Truncated Wigner simulation show a remarkable agreement with our theory.
@@
The results of Secs. V-VI can be used to describe simultaneous symmetry breaking in both spontaneous and conventional Floquet states. Furthermore, the $(t,\phi)$ formalism and the Floquet thermodynamics can be applied to an arbitrary many-body system, where they establish a fundamental distinction between both states at the thermodynamical level, leading to an important conclusion: spontaneous Floquet states do not experience heating as they conserve energy, in contrast to conventional Floquet systems D’Alessio and Rigol (2014); Lazarides et al. (2014). In the specific case of the CES state, this survives for thermodynamically long times which scale linearly with the system size de Nova and Sols (2022). This can be compared with the prethermal regime of a driven Floquet system, whose lifetime scales exponentially with the applied frequency Pizzi et al. (2021); Ye et al. (2021); Kyprianidis et al. (2021).

### v4

@@ 7 Time supersolids: CES state
Figure 5: (a) Normalized density-density correlation function G(x,x,t)G(x,-x,t) for $x=10$ (blue), $x=20$ (red), and $x=30$ (green) for the simulation in Fig. 4. Black line is a quadratic polynomial fit to the amplitude of the oscillations, $A(t)=at^{2}+bt+c$. (b) Zoom of (a). (c)-(d) Same as upper row but now replacing quantum fluctuations by particle-number fluctuations. Specifically, we take as initial condition in the Truncated Wigner method $\Psi(x,0)=\sqrt{1+\delta n}\,e^{ivx}$, with $\delta n$ a Gaussian random variable with zero mean and $\sqrt{\braket{\delta n^{2}}}=0.001$.
We now go beyond mean-field and explicitly compute the dynamics of the quantum fluctuations by using the Truncated Wigner method [60, 61]; all technical details behind the numerical simulations and the specific implementation of the Truncated Wigner method can be found in Refs. [44, 43, 98]. As initial quantum state, we take the $T=0$ ground state in the comoving frame of the condensate. Regarding the observables of interest, we focus on computing the normalized density-density correlation function
@@
with $\delta\hat{n}(x,t)\equiv\hat{n}(x,t)-n(x,t)$ the fluctuations around the ensemble-averaged density $n(x,t)\equiv\braket{\hat{n}(x,t)}$, and $\xi_{0}$ the healing length associated to the initial density $n_{0}$; we have momentarily restored dimensions in this expression for the sake of clarity. In the laboratory, density-density correlations can be measured through in situ imaging after averaging over ensembles of repetitions of the experiment [99, 100, 101].
After some transient, and once in the CES state, the correlations exhibit a seemingly periodic behavior, as shown in Fig. 4, where sharp features arise due to the synchronized emission of shock waves/solitons into the upstream/downstream region. Further insight is obtained in upper row of Fig. 5, where we plot the time dependence of the correlation function G(x,x,t)G(x,-x,t) between symmetric upstream and downstream points for different values of $x$. We observe that the correlations are indeed quasi-periodic, displaying periodic oscillations whose amplitude grows in time.
@@
Figure 6: Normalized density-density correlation function G(x,x,t)G(x,-x,t) for $x=10$ (left column), $x=20$ (center column), and $x=30$ (right column). Solid blue is the numerical result from a Truncated Wigner simulation and dashed red is the contribution from the temporal FNG mode, $G(x,x^{\prime},t)=A(t)r_{t}(x,t)r_{t}(x^{\prime},t)$, where $A(t)$ is numerically fitted as in Fig. 5 and $r_{t}(x,t)=\partial_{t}n_{0}(x,t)$ is computed from the time derivative of the mean-field CES density. Upper row: $Z=1$ and $v=0.8$. Lower row: $Z=1$ and $v=0.7$.
This behavior can be understood using the results of Sec. 6. After expanding the field operator around the CES wavefunction, we find that the density fluctuations read at linear order in the field fluctuations $\hat{\varphi}(x,t)$ as
@@
$\delta\hat{n}(x,t)=\hat{X}^{t}(t)r_{t}(x,t)+\hat{P}^{N}r_{N}(x,t)+\hat{P}^{F}r_{F}(x,t)\\ +\sum_{n}\hat{\gamma}_{n}r_{n}(x,t)e^{-i\varepsilon_{n}t}+\hat{\gamma}^{\dagger}_{n}r^{*}_{n}(x,t)e^{i\varepsilon_{n}t},$ (180)
$r(x,t)=-iz^{\dagger}_{\theta}(x,t)\sigma_{z}z(x,t)$ being the density wavefunction associated to the spinor $z$. In the case of the Goldstone-Gibbs modes, the density wavefunction is simply obtained by
$r_{a}(x,\phi)=\partial_{a}|\Psi_{0}(x,\phi)|^{2}\equiv\partial_{a}n_{0}(x,\phi),$ (181)
where $n_{0}(x,\phi)$ is the mean-field CES density, which is approximately equal to the ensemble-averaged density after neglecting higher-order corrections in the field fluctuations, $n(x,t)\simeq n_{0}(x,t)=n_{0}(x,\phi_{0}+\omega t)$.
@@
$G(x,x^{\prime},t)$ $\simeq A(t)r_{t}(x,t)r_{t}(x^{\prime},t),$ (182)
$A(t)$ $\equiv\frac{\braket{(\hat{X}^{t}(t))^{2}}}{n_{0}\xi^{-1}_{0}}.$
@@
This is a quasi-periodic time dependence, where $r_{t}(x,t)r_{t}(x^{\prime},t)$ is a periodic term as $r_{t}(x,t)=\partial_{t}n_{0}(x,t)$, while the global amplitude $A(t)$ encodes the non-periodic part. Specifically, due to the ballistic evolution of $\hat{X}^{t}(t)$, we expect a quadratic polynomial dependence, $A(t)=at^{2}+bt+c$. This fits well the observed growth, black lines in upper row of Fig. 5. A more detailed comparison is presented in Fig. 6, where we depict together the numerical results for $G(x,x^{\prime},t)$ from the Truncated Wigner method (solid blue), and the theoretical prediction from Eq. (182), combining the previous quadratic fit of $A(t)$ with a computation of $r_{t}(x,t)$ from the time derivative of the mean-field density (dashed red). A remarkable agreement is observed, confirming the ballistic motion of the quantum amplitude of the temporal FNG mode as the mechanism responsible for the quasi-periodic correlations.
@@
We note that the coefficients $a,b,c$ determining the amplitude $A(t)$ are given in terms of expectation values which are quadratic in $\hat{X}^{t},\hat{P}^{N},\hat{P}^{F}$, evaluated in the initial quantum state. Thus, our specific time-dependent scheme eliminates any possible ambiguity in the definition of the quantum state of the system [102]. Nevertheless, fluctuations of the temporal FNG can be also induced by run-to-run experimental variations such as shot noise in the initial particle number. This is shown in lower row of Fig. 5, where we display the results of solely including fluctuations in the particle number of the initial condition. The observed correlation patterns are the same as in the purely quantum case (upper row), further revealing their monomode origin since particle-number fluctuations are structureless [98].
@@
Based on these results, we propose an experimental configuration to observe the temporal FNG mode involving standard techniques in analogue gravity setups [99, 100, 101], schematically depicted in Fig. 7. We consider an elongated quasi-1D condensate confined by a boxlike potential [103, 104, 105, 106, 107], whose density is essentially homogeneous far from the edges. At $t=0$, a localized obstacle (indicated by the shaded area) is swept with velocity $v$ along the bulk of the condensate. By Galilean invariance, this is equivalent to launching the condensate against the obstacle with the same velocity, achieving a configuration similar to that in Fig. 3a. For values of $v$ above the critical velocity, a CES state will be reached at long times. Another possibility is to confine the condensate in a long ring [108] and rotate a localized potential. In both cases, at late times once in the CES regime, when the temporal FNG mode dominates the Bogoliubov dynamics, high-resolution imaging can be then used to measure the time evolution of the density profile $n(x,t)$ and the density-density correlations $G(x,x^{\prime},t)$, from where the FNG mode can be extracted as in Figs. 4-6.
@@
We note that the protocol used here to reach the CES state is fully deterministic at the mean-field level, which imposes the value of the global time origin $t_{0}$ of the spontaneous Floquet state. Thus, even though there is a temporal FNG mode, technically speaking the symmetry breaking of time-translation invariance is not spontaneous. The reason is that our scheme fixes a $t=0$ origin in the GP equation. Regardless, at late times, due to interactions, the system enters the CES self-oscillating regime where it forgets about the initial condition and exhibits all the features of a spontaneous Floquet state, including the temporal FNG mode. We can qualitatively understand this behavior by drawing an analogy with a conventional spatial crystal growing in the $z$-direction on top of a rigid substrate, which fixes a $z=0$ plane that breaks continuous spatial translation symmetry. In the bulk of the grown crystal, however, the system becomes insensitive to the boundary and recovers the usual properties of a crystalline structure.
Figure 7: Schematic representation of the proposed experimental setup to observe the FNG mode. A localized obstacle (indicated by the shaded area) is swept with velocity $v$ along the homogeneous bulk of a condensate confined by a boxlike potential (whose initial density profile is given by solid blue line). By Galilean invariance, this is equivalent to launching the condensate over the obstacle with the same velocity. When $v$ is above the critical velocity, the CES state is achieved at long times; then its density correlations can be characterized through high-resolution imaging.
@@
Nonetheless, the phase locking of $\phi_{0}$ actually helps to measure the FNG mode since, if $\phi_{0}$ were spontaneously chosen in each realization of the experiment, then the ensemble-averaged value of $G(x,x^{\prime},t)$ would represent an angular average over $\phi_{0}\in[0,2\pi)$, therefore suppressing its time dependence. In that case, one would need to resort to the complete out-of-time correlation function
@@
which requires the use of non-destructive imaging in order to measure the density at two consecutive times $t>t^{\prime}$, as in Ref. [21]. This caveat is not needed for the measurement of correlations under the usual $U(1)$ symmetry breaking, where each realization of the experiment is insensitive to the global condensate phase because observables always commute with the total particle number, the generator of global phase transformations.
8 Discussion and perspectives
@@
We proceed to critically discuss the results of this work and how they inscribe within the literature, outlining in the process future avenues. We have examined in Sec. 3 how the general results from Goldstone theorem applied to variational descriptions, Sec. 2, can be translated to many-body systems when several symmetries are simultaneously broken by the quantum state, focusing on the specific case of a BEC close to zero temperature for illustrative purposes. We have found that a rather complete description of the problem is provided in terms of the generalized Gibbs ensemble, which treats all broken symmetries on equal grounds. The Lagrange multipliers associated to each conserved charge behave as generalized velocities along the orbits of the symmetry transformations, extending the well-known result that the chemical potential is the velocity of the phase. We have also proven fundamental results in Thermodynamics for the symmetry-broken state, such as the first principle or the Gibbs-Duhem relation.
@@
The Nambu-Goldstone modes associated to each spontaneously broken symmetry emerge as zero-energy modes of the excitation spectrum. Each NG mode is paired with a Gibbs mode, describing the fluctuations of the quantum state with respect to the associated conserved charge. These fluctuations arise because a symmetry-broken state cannot be eigenstate of the corresponding charge operators. The orthogonalization of the Goldstone-Gibbs modes involves the Berry-Gibbs connection, a Berry connection whose variables are the conserved charges, which are not intrinsic parameters of the Hamiltonian. In contrast to the standard case, the resulting Berry-Gibbs curvature is not invariant under generalized gauge transformations, expanding the usual notion of phase transformations to include more general symmetry transformations. When extended to the complete manifold that includes the continuous parameters of the spontaneously broken symmetries, the Berry-Gibbs curvature gives rise to a symplectic structure that allows to quantize the amplitudes of the Nambu-Goldstone/Gibbs modes as conjugate coordinate/momentum operators. This analogy is translated to the dynamics, where the momenta are time-independent as they represent the quantum fluctuations of the corresponding conserved charges, while the coordinates display a ballistic motion with velocities that depend linearly on the momenta. The physical intuition behind this dynamics is very simple: since there is no restoring force along the orbits generated by the broken-symmetry transformations, fluctuations in those directions, spanned by the NG modes, lead to unbounded motion with constant velocity. Finally, we show that the Goldstone-Gibbs sector can be diagonalized, leading to a new set of independent NG modes. Remarkably, this diagonalization only involves thermodynamic considerations, and describes how the original NG modes hybridize within the different branches of the spectrum in the low-energy limit.
@@
The coordinate-momentum analogy between the Nambu-Goldstone and Gibbs amplitudes was originally established in Ref. [55] in the context of the $U(1)$-symmetry breaking and the phase diffusion of a condensate. This approach was followed in Ref. [56] to also quantize the position of a soliton and study its quantum diffusion, inspiring a whole line of research on quantum solitons [109, 110, 111, 112, 113]. A study of the quantization of arbitrary NG modes in condensates, including internal symmetries such as spin, was provided in Ref. [114]. The present work further extends those results by developing a general and systematic framework with the help of the generalized Gibbs ensemble and the Berry-Gibbs connection, which in addition allows to understand the physics of the Nambu-Goldstone and Gibbs modes at a more fundamental level. On the other hand, the emergence of a geometrical description and a symplectic structure in variational formulations is well-known in the literature and can be traced back to Ref. [115] (see Ref. [116] for a more recent review), where the manifold is generated there by the variational parameters of the ansatz and their velocities, leading to a dequantization of the problem. In contrast, the manifold variables in our description are the continuous parameters of the spontaneously broken symmetries and their associated charges, instead of variational parameters, and the symplectic structure conversely leads to a quantization of the Goldstone-Gibbs amplitudes. To the best of our knowledge, the thermodynamical and geometrical framework developed in Sec. 3 for the description of spontaneous symmetry breaking and the quantization of the Goldstone-Gibbs modes is original from this work and is not present in the literature.
@@
As a particular application of this formalism, we have studied in Sec. 4 the case of a cnoidal wave in a superfluid, which spontaneously breaks $U(1)$ and spatial translation symmetry, extending the work of Refs. [56, 57] into the generalized Gibbs ensemble. We have explicitly checked the consistency and validity of the theoretical predictions of Sec. 3.
@@
The results of Secs. 3, 4 are thus of interest for the study of states which simultaneously break several continuous symmetries, such as supersolids, quantum droplets, spinor and rotating condensates, or quantum solitons. In particular, they find immediate application in systems described by extended nonlinear equations beyond the usual Gross-Pitaevskii description, such as the Lee-Huang-Yang correction [117, 118, 119] or non-polynomial Gross-Pitaevskii equations [120, 100]. Our results can also be easily translated to magnonic condensates [121], nonlinear optical fibers [122], and quantum fluids of light [123], as these obey similar nonlinear equations of motion. Indeed, the generality of the formalism, based on a Lagrangian description, allows its adaptation to arbitrary many-body systems, including fermions. For instance, within the $\nu=0$ quantum Hall phase diagram in graphene, states with two spontaneously broken symmetries were predicted within a Hartree-Fock description [124], of which some experimental evidence was recently reported [7]. Indeed, Ref. [124] established the correspondence between the BdG and TDHFA equations, and translated Goldstone theorem into the specific framework of the Hartree-Fock equations. Therefore, the BdG/TDHFA correspondence provides a direct pipeline through which our results can be exported.
@@
Another interesting perspective is provided by the geometry of the extended Berry-Gibbs connection. In this work, we have assumed that all the generators of the spontaneously broken symmetries commute between themselves. This leads, after a generalized gauge transformation, to a “flat” symplectic curvature $F_{ab}=-\Omega_{ab}$. Actually, a generalized gauge transformation is nothing else but a change of coordinates in the extended Berry-Gibbs manifold. However, if the generators of the spontaneously broken symmetries form a non-abelian Lie algebra, this no longer holds, opening the door for deeper geometrical implications, including non-trivial topological aspects. The exploration of this promising avenue is left for future work.
@@
The formalism developed for stationary states in Secs. 2, 3 is adapted to Floquet states in Secs. 5, 6. Specifically, from the general results of Sec. 2, we prove that the usual Nambu-Goldstone modes with zero energy are now translated into Floquet-Nambu-Goldstone modes with zero quasi-energy. Moreover, since spontaneous Floquet states break continuous time-translation symmetry, they possess a genuine temporal FNG mode. This is a distinctive signature of a spontaneous Floquet state, completely absent in conventional Floquet systems.
@@
In order to correctly describe the dynamics of the FNG modes of a spontaneous Floquet state, we develop the $(t,\phi)$ formalism, which goes beyond the usual $(t,t^{\prime})$ formalism. In combination with the generalized Gibbs ensemble, we provide a thermodynamical description of Floquet states analogous to that of stationary states, which we refer to as Floquet thermodynamics. This is possible because spontaneous Floquet states also have a well-defined energy, which enables the existence of a conserved Floquet charge. The use of the phase $\phi$ instead of the periodic time $t^{\prime}$ is essential to both identify the thermodynamical role of the Floquet charge $Q_{\phi}=F$ and the dynamical role of its Lagrange multiplier $\lambda^{\phi}=\omega$. The Floquet charge is another hallmark of a spontaneous Floquet state, which adds to the presence of a temporal FNG mode. In fact, both features are the sides of the same coin, since they arise, via Noether and Goldstone theorems, from the continuous time-translation symmetry of the underlying time-independent Hamiltonian.
@@
Finally, we apply our formalism to a specific realization of spontaneous Floquet state, the CES state, which spontaneously breaks $U(1)$ and continuous time-translation symmetry, representing a time supersolid. We propose an experimental scheme, based on standard analogue gravity techniques [99, 100, 101], to observe the temporal FNG mode, which does couple to the density in contrast to the $U(1)$ FNG mode. In particular, we prove that the density-density correlations are dominated at long times by the ballistic growth of the temporal FNG mode, from where it can be extracted. Numerical results from a Truncated Wigner simulation show a remarkable agreement with our theory.
@@
The results of Secs. 5, 6 can be used to describe simultaneous symmetry breaking in both spontaneous and conventional Floquet states. Furthermore, the $(t,\phi)$ formalism and the Floquet thermodynamics can be applied to an arbitrary many-body system, where they establish a fundamental distinction between both states at the thermodynamical level. In the specific case of the CES state, this survives for thermodynamically long times which scale linearly with the system size [43]. This can be compared with the prethermal regime of a driven Floquet system, whose lifetime scales exponentially with the applied frequency [125, 126, 23]. From a broader perspective, the description of general quantum systems in terms of traditional thermodynamic tools is an active topic of research [127, 128, 129, 130].